A 375 g cup of coffee cools as it sits on the table. The coffee was 97C and cools to 25C. *




1
Expert's answer
2020-11-26T05:08:07-0500

To find the heat lost, we use the formula,

H = mc∆T

m = 375g

∆T (change in temperature) = 97 - 25 = 72°C

c (specific heat capacity of coffee) = 1.4J/g°C


"\therefore" H = 375 × 1.4 × 72 = 37800J = 37.8kJ


Therefore, 37.8kJ of heat was lost as it cools.